OpenStudy (meilendurcer):

The length of the hypotenuse of a right triangle is 13 m. The length of one leg is 5 cm. Find the length of the other leg.

OpenStudy (agent0smith):

 Use pythagoras \[\large 13^2 = a^2 + 5^2\] Now try to solve for a.

OpenStudy (meilendurcer):

8?

OpenStudy (agent0smith):

You can't just drop the square roots from all of them. \[\large 13^2 = a^2 + 5^2 \] \[\large 169 = a^2 + 25\]